OpenStudy (anonymous):

A 10−gallon container of water costs (y + 2) dollars while a 5−gallon container of water costs (y − 2) dollars. If the price per gallon of water is the same for both containers, the equation below can be used to solve for y. What is the value, in dollars, of y? (Do not put a $ in your answer.)

OpenStudy (jackellyn):

y+2/10 = y-2/5 cross multiply 5y+10 =10y-20 30 = 5y y=6